# Investigating LLM Capabilities on Long Context Comprehension for Medical Question Answering

### v1

@@ Abstract
This study is the first to investigate LLM comprehension capabilities over long-context (LC) medical QA of clinical relevance. Our comprehensive assessment spans a range of content-inclusion settings based on their relevance, LLM models of varying capabilities and datasets across task formulations, revealing insights on model size effects, limitations, underlying memorization issues and the benefits of reasoning models. Importantly, we examine the effect of RAG on medical LC comprehension, uncover best settings in single versus multi-document reasoning datasets and showcase RAG strategies for improvements over LC. We shed light into some of the evaluation aspects using a multi-faceted approach. Our qualitative and error analyses address open questions on when RAG is beneficial over LC, revealing common failure cases.
@@
Feras AlMannaa1, Talia Tseriotou2, Jenny Chim2, Maria Liakata2,3 1Istanbul Aydın University 2Queen Mary University of London 3The Alan Turing Institute {ferasmhdaymanalmanna}@stu.aydin.edu.tr,{t.tseriotou,m.liakata}@qmul.ac.uk
@@ 1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Dubey et al., 2024; Yang et al., 2024) perform impressively on medical tasks (Borgeaud et al., 2021; Nori et al., 2023), achieving superhuman scores on USMLE-style exam question-answering (QA) (Chen et al., 2023; Tang et al., 2023; Pal and Sankarasubbu, 2024; Singhal et al., 2025). Yet multiple-choice QA (MCQA) results are not indicative of performance in more complex tasks Arias-Duart et al. (2025), including open-ended medical QA (Sandeep Nachane et al., 2024) and long-form generation. Furthermore, board exam questions rely primarily on textbook knowledge, likely exposed during pre-training (Chen et al., 2025).
Beyond MCQA, researchers have constructed datasets based on authentic de-identified electronic health records (EHRs). While these benchmarks still face challenges regarding data diversity, ecological, external, and construct validity Wornow et al. (2023), they provide crucial signals on how well models can address complexities in real data, including domain-specific vocabulary, heterogeneous document types, multi-document reasoning, data noise, linguistic diversity, long contexts, and long-range dependencies.
@@
To date most EHR-based benchmarks involve tasks that only require single-document contexts (Pampari et al., 2018; Yue et al., 2020). In practice, there is no guarantee that the necessary information (e.g. to perform a diagnosis) can be found in a single document known a priori; what is needed is comprehension and reasoning over largely noisy and at best complimentary records. EHR-Note-QA (Kweon et al., 2024) is the first to include questions requiring information across records including both multiple-choice (MC) and open-ended questions. There is limited work on the effect of long-range dependencies within biomedical machine reading comprehension (MRC) Vatsal and Singh (2024) and this tends to focus on shorter texts ($\sim$1.5-4K tokens) and biomedical articles and span-based QA rather than complex EHRs. Work that addresses MRC in other domains has utilised methods such as vector compression (Chevalier et al., 2023), memory-based models (Xiao et al., 2024) and context window chunking or sliding (Xiao et al., 2023; Han et al., 2023) to process long context.
Retrieval Augmented Generation (RAG) (Lewis et al., 2021) forms a cost efficient alternative that mitigates risks around information loss or utilizing outdated or incomplete knowledge by selecting relevant context on demand rather than trying to process the entire context. What is preferable is subject to debate, with each showing benefits on different tasks and settings (Li et al., 2024): RAG has been reported to enhance LLM performance compared to long-context prompt compression for QA (Zhang et al., 2024) outperforming SOTA LLMs for both 32K and 128K benchmarks (Xu et al., 2024; Bai et al., 2024). By contrast, others showed that RAG performance peaks at 32K context length for large LLMs (Leng et al., 2024).
Here we investigate how LLMs handle diverse input patient record combinations, using EHR-based human-verified datasets in a variety of long context and RAG settings. Fig. 1 shows a diagram of our study setup. We make the following contributions:
  • We are the first to assess how LLMs of different underlying capabilities and sizes perform on long-context medical QA given longitudinal EHR patient notes. We cover three task formulations: 1) MC , 2) Extractive and 3) Open-ended generative QA. (§3.13.2)
  • We perform comprehensive experiments evaluating the effect of different levels of granularity in the input context while also testing for EHR memorisation (§4.2).
  • We examine the effect of RAG across task formulations and context sizes, showing that given the right setting RAG has superior performance than LC across tasks (§5).
  • We examine the distinct challenges of open-ended QA when reasoning over a single document versus multiple documents (§5).
  • Our qualitative analyses shed light on some of the challenges that LLMs currently face in reasoning over long EHRs (§5.1).
@@ 2.1 Leveraging Long Context
While the increasing long-context window capabilities of some LLMs, i.e. GPT-4o (Hurst et al., 2024) (128K), Claude 3 (Anthropic, 2023) (200K), Gemini 1.5 (Team et al., 2024) (1 million), boost their long-context performance they still struggle in general purpose open-form QA (beyond short passage retrieval) Zhang et al. (2024); Chen et al. (2025). Despite positional embeddings techniques like RoPE (Su et al., 2024), YaRN (Peng et al., 2023) and Position Interpolation (Chen et al., 2023) allowing context extrapolation, QA task performance is low even for 32K context (Chen et al., 2025). Furthermore, the performance gap between open-source and longer-context close-source models (Li et al., 2024) poses questions around the potential avenues for patient-centric long-dependency data where privacy matters.
Benchmarking and assessment of long-context performance of LLMs (Dong et al., 2023; Bai et al., 2024; Liu et al., 2023; Zhang et al., 2024; Hsieh et al., 2024), shows that apart from performance drops with increasing token length LLMs are particularly challenged by long-range dependency QA tasks (Li et al., 2024) that potentially require reasoning. Fan et al. (2024) show notable LLM performance deterioration in medical QA tasks with increased context length up to 200K tokens, with open-source LLMs struggling to produce output given larger contexts. Although long and longitudinal context is prominent in real-world patient and medical data, there is little work investigating LLM capabilities in such settings.
@@ 2.2 Retrieval-augmented Generation (RAG)
RAG has been leveraged in a prompt-based plug-in manner to address black-box long-context challenges (Yu et al., 2023). Current retrieval strategies are primarily: chunk-based (Izacard et al., 2021), index-based (Liu, 2022) and summarization-based (Sarthi et al., 2024). LLM-based large-scale retrievers for chunk-based approaches like E5-Mistral-7b (Wang et al., 2022) and BGE M3 (Chen et al., 2024) give notable performance improvement. Furthermore, approaches like Self-RAG (Asai et al., 2023) introduce on-demand retrieval followed by generation and self-reflection outperforming ChatGPT in open-domain QA.
Domain-specific retrievers strongly benefit biomedical retrieval. MedCPT (Jin et al., 2023), a first of its kind contrastive biomedical dense semantic retriever and re-ranker, achieved top performance on 6 biomedical tasks surpassing LLM-based retrievers. BMRETRIEVER (Xu et al., 2024), a biomedical dense retriever trained both contrastively and instruction fine-tuned on synthetic data, shows better performance than MedCPT on most downstream tasks. The Med-RAG (Zhao et al., 2025) method provides better diagnostics over patient EHRs, being enhanced by medical knowledge graph-elicited reasoning. Finally, the Self-BioRAG (Jeong et al., 2024) is a framework that retrieves relevant medical documents using MedCPT and generates through self-reflection on rationales using a domain-specific LLM. Yet Fan et al. (2025) underscore the challenges in retrieval for the medical domain: despite strong medical retrieval capabilities across large corpora, models struggle with specialized medical content such as EHRs. Myers et al. (2025) further demonstrated variability in retrieval across EHR datasets. There is need for work that assesses retrieval performance on patient-oriented QA over long-context.
@@ 3.1 Task Selection
To uncover the limitations of LLMs in medical QA given complex patient-oriented clinical scenarios, we focus on content-comprehension settings especially ones where patient context is not supplementary but crucial to complete a task (e.g. obtain diagnoses), thus emphasizing long temporal patient context spanning across time and documents.
@@
Aiming for a comprehensive approach we evaluate datasets covering three task formulations: 1) MCQA, 2) Extractive, and 3) Open-ended QA. We further assess each dataset across four settings with varying note relevance.
3.2 Dataset Selection
@@
Datasets were included in our study based on the following criteria (datasets comparison in Table LABEL:tab:dataset_landscape):
@@
Going beyond Open-domain MCQA: Widely used medical MC benchmarks i.e. MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), PubMedQA (Jin et al., 2019), MMLU (Hendrycks et al., 2021) evaluate option selection instead of grounded synthesis and risk pretraining leakage from exams and textbooks, which can inflate scores, while importantly providing little signal towards noisy clinical settings.
@@
Long-context Documents: To address LLM capabilities in handling long context that reflects real-world settings we seek benchmarks that extend well-beyond 8K. To avoid a simplified needle in the haystack setting we aim for such content to be relevant rather than distracting.
Expert QA Annotations: We selected expert annotated QA pairs grounded on multiple EHR documents to avoid limitations that hinder clinical reliability such as non-expert curation (Fan et al., 2024) or synthetic data (Adams et al., 2024). Patient-centric Longitudinal Content: To study performance on clinically-relevant QA with multi-note reasoning, we prioritize EHR datasets derived from MIMIC-III and MIMIC-IV (Johnson et al., 2016, 2023). These corpora offer standardized de-identified notes across document types (e.g., discharge summaries, radiology reports) and enable consistent retrieval granularity.
3.3 RAG Strategy
@@
RAG strategies can be distinguished into sparse and dense. Sparse retrievers such as BM25 (Robertson et al., 2009) and Splade (Formal et al., 2021) use overlapping terms to match queries with snippets, while dense retrievers encode them into embeddings and match them based on semantic similarity. Recently, using LLM-based embeddings such as E5-Mistral-7B-Instruct (Wang et al., 2023) and Qwen3 Embedding (Zhang et al., 2025) for dense retrieval has been standard practice.
@@
Research on biomedical MCQA (Wang et al., 2024) and biomedical document retrieval (Luo et al., 2022) has shown that hybrid approaches of sparse and dense retrieval outperform single component counterparts. Further, re-ranking of retrieved snippets also exhibits superior performance in biomedical applications (Jin et al., 2023; Wang et al., 2024; Sohn et al., 2025). Additionally, non-hybrid experiments on different EHR subsets of varying granularity (Fan et al., 2025) show that there is no clear performance winner between sparse and dense retrievers. Therefore, we use a hybrid approach of a sparse and denser retriever with a re-ranker.
Our hybrid approach is presented in Figure 2. Clinical notes are segmented into 512-token chunks, which form our document collections $D=\{d_{1},d_{2},\dotsc,d_{n}\}$. The question respectively forms the query $Q$.
@@
For sparse retrieval we employ a lexical retriever as a ranking function of each chunk $d_{j}$ that extends TF-IDF by considering term frequency saturation and document length  [^1] Here terms document and chunk are used interchangeably.:
@@
For the dense retriever we use the same Encoder, $E$ to encode the query and each of the chunks and capture semantic similarity of pairs using the cosine similarity:
@@
Top $k$ chunks are retrieved with each retriever and then combined through Reciprocal Rank Fusion (RRF) that takes into account the rank of each document:
@@
, where $k_{RRF}$ is a smoothing constant.
Finally, we employ Late Interaction Reranking, a measure of how well every query token is semantically supported by at least one document token, to produce the final top $k$ most contextually relevant set.
@@
Document Ordering: We leverage an order-preserving RAG that sorts retrieved chunks temporally rather than by retrieval score, due to its superior long-context performance (Yu et al., 2024) regardless of the number of chunks.
Note Inclusion: We evaluated two different strategies: direct chunk inclusion, and parent note inclusion (§4.3).
@@ 3.4 Evaluation Methodology
Evaluating medical QA systems has traditionally relied on automatic metrics and multiple-choice benchmarks Jin et al. (2019, 2021); Pal et al. (2022), which probe memorized knowledge but underrepresent the complexity and factual rigor of clinical reasoning, long-form answers, and EHR-grounded context. To provide clinically meaningful assessment, we adopt a multi-dimensional, reference-based scheme that complements lexical overlap with embedding-based semantic similarity and domain-adapted Natural Language Inference (NLI) for factual consistency and precision. In line with emerging clinical evaluation practices (e.g., MEDIC (Kanithi et al., 2024)) and the rise of LLM-as-a-judge, we employ a calibrated rubric to capture Correctness, Completeness, and Faithfulness at scale, and we prioritize widely available, reproducible metrics while cross-checking signals for robustness. Given that gold standard answers are available, we employ reference-based evaluation:
METEOR Banerjee and Lavie (2005) is used to capture surface-level and lexical similarity.
BERTScore Zhang et al. (2020) computed with Clinical BioBERT embeddings Alsentzer et al. (2019) assesses the semantic similarity.
NLI scores capture the logical relationship between reference and candidate answers. We use a domain-adapted NLI model Deka et al. (2023) with reference answers as premise and candidate answers as hypotheses. We then measure Factual Consistency using the probability of non-contradiction Song et al. (2024) and measure Factual Precision using the entailment probability.
LLM-as-a-judge was employed with a 5-point scale rubric on three aspects, using Qwen-2.5-32B-Instruct as the evaluator model [^2] We selected Qwen for stability and agreement after comparing against Selene-8B (Alexandru et al., 2025) and Prometheus-8x7B v2.0 (Kim et al., 2024) in a small pilot (see Appendix LABEL:sec:judgecomps).. Correctness captures factual consistency with respect to the gold answer, penalizing contradictions while allowing compatible additional information. Completeness captures recall by assessing the predicted answer’s coverage of information present in the reference. Faithfulness captures precision by assessing whether the predicted answer only contains information that is supported by or derivable from the reference.
We report Accuracy on the MC task.

### v2

@@ Abstract
This study is the first to investigate LLM comprehension capabilities over long-context (LC), clinically relevant medical Question Answering (QA) beyond MCQA. Our comprehensive approach considers a range of settings based on content inclusion of varying size and relevance, LLM models of different capabilities and a variety of datasets across task formulations. We reveal insights on model size effects and their limitations, underlying memorization issues and the benefits of reasoning models, while demonstrating the value and challenges of leveraging the full long patient’s context. Importantly, we examine the effect of Retrieval Augmented Generation (RAG) on medical LC comprehension, showcasing best settings in single versus multi-document QA datasets. We shed light into some of the evaluation aspects using a multi-faceted approach uncovering common metric challenges. Our quantitative analysis reveals challenging cases where RAG excels while still showing limitations in cases requiring temporal reasoning.
@@
Feras AlMannaa1∗, Talia Tseriotou2∗, Jenny Chim2, Maria Liakata2,3 1Istanbul Aydın University 2Queen Mary University of London 3The Alan Turing Institute {ferasmhdaymanalmanna}@stu.aydin.edu.tr,{t.tseriotou,m.liakata}@qmul.ac.uk
* Indicates equal contribution.
@@ 1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Grattafiori et al., 2024; Yang et al., 2024) perform impressively on medical tasks (Borgeaud et al., 2022; Nori et al., 2023), achieving superhuman scores on United States Medical Licensing Examination (USMLE)-style exam question-answering (QA) (Chen et al., 2023b; Tang et al., 2024; Pal and Sankarasubbu, 2024; Singhal et al., 2025). Yet evaluation is mostly based on multiple-choice QA (MCQA) (Liévin et al., 2024; Xiong et al., 2024; Chen et al., 2024b; Singhal et al., 2025) which doesn’t reflect performance in complex tasks Arias-Duart et al. (2025), such as open-ended medical QA (Sandeep Nachane et al., 2024). Furthermore, medical board exams designed to assess professional knowledge and decision-making rely primarily on textbook knowledge which is likely available during pre-training (Chen et al., 2025a). Assessing LLM capabilities on context from expert-curated electronic health records (EHRs) could provide crucial signals on how well models address complexities in real data, including domain-specific vocabulary, heterogeneous document types, multi-document reasoning, data noise, linguistic diversity, long contexts, and long-range dependencies Wornow et al. (2023). However, research on open-form QA pertaining to EHRs focuses on creating datasets (Yang et al., 2022) or EHR-specialized models (Fleming et al., 2024), rather than investigating LLM performance in real-world QA.
@@
Moreover, most EHR-based benchmarks consider single-document contexts (Pampari et al., 2018; Yue et al., 2021), while in practice there is no guarantee that the necessary information will be included within a specific patient document. Reasoning over multiple long, temporally-dependent documents further poses questions around long-context (LC) limitations and evidence positioning sensitivity (Liu et al., 2024; Xiao et al., 2023), also observed in medically focused studies (Adams et al., 2025). Work on the effect of long-range dependencies within medical machine reading comprehension (MRC) is limited Vatsal and Singh (2024), focusing on shorter texts ($\sim$1.5-4K tokens) and span-based QA rather than complex EHRs. Work on LC medical QA involves either synthetic data, is MCQA-based and still relatively short (up to 6k) (Adams et al., 2025)), or small in sample size and not human-validated (Fan et al., 2025a). We address this gap by investigating LLM capabilities to answer questions given longitudinal patient-centric EHRs, where QA pairs are human-validated.
While recent LLM releases push the frontiers of context size, performance does not scale with increased context (Levy et al., 2024; Modarressi et al., 2025). Retrieval Augmented Generation (RAG) (Lewis et al., 2020) often forms a cost efficient alternative by selecting relevant context on demand. Whether LC or RAG is preferable is subject to debate, with each showing benefits for different tasks and settings (Li et al., 2024b). RAG has been reported to enhance LLM performance compared to LC prompt compression for QA (Zhang et al., 2024a), outperforming SOTA LLMs on both 32K and 128K benchmarks (Xu et al., 2024a; Bai et al., 2025). By contrast, others showed that RAG performance peaks at 32K context length for large LLMs (Leng et al., 2024). While work on medical QA reports promising gains when retrieval is carefully designed (Xiong et al., 2024), LC vs RAG superiority in long-form medical QA remains an open question. We make the following contributions:
  • We are the first to assess how LLMs of different underlying capabilities and sizes perform on long-context medical QA given longitudinal EHR patient notes, covering three task formulations: 1) MCQA , 2) Extractive and 3) Open-ended generative QA (§4.14.3) and a variety of LC and RAG settings, as depicted in Fig. 1.
  • We perform comprehensive experiments evaluating different note inclusion strategies in the input context, including note memorization (§4.2).
  • We carefully design a hybrid RAG pipeline examining its performance over Full-context (FC) across task formulations and context sizes, showing its superiority across tasks (§5.15.2).
  • Our experiments and multi-faceted evaluation methodology allow for a direct comparison between long-form reasoning when the answer is part of a single versus multiple documents (§5.1).
  • Our quantitative and qualitative analyses uncover some of the LLM and metrics challenges in reasoning over long EHRs, while demonstrating such cases where RAG excels (§5.25.3).
@@ 2.1 Leveraging Long Context
While increasing long-context window capabilities of LLMs, e.g. GPT-4o (Hurst et al., 2024) (128K), Claude 3 (Anthropic, 2023) (200K), Gemini 1.5 (Team et al., 2024a) (1 million), boost their long-context performance they still struggle in general purpose open-form QA (beyond short passage retrieval) Zhang et al. (2024b); Chen et al. (2025b). Despite positional embedding techniques like RoPE (Su et al., 2024), YaRN (Peng et al., 2023) and Position Interpolation (Chen et al., 2023a) allowing context extrapolation, QA task performance is low even for 32K contexts (Chen et al., 2025b). Furthermore, the performance gap between open-source and longer-context close-source models (Li et al., 2024a) poses questions around the potential avenues for patient-centric long-dependency data where privacy matters.
Assessment of long-context LLM performance (Dong et al., 2024; Bai et al., 2024; Liu et al., 2024; Zhang et al., 2024b; Hsieh et al., 2024), shows that apart from performance drops with increasing token length, LLMs are particularly challenged by long-range dependencies (Li et al., 2024a) that potentially require reasoning. Fan et al. (2025a) show notable LLM performance drop in medical QA tasks with contexts up to 200K tokens, with open-source LLMs struggling to produce output given larger contexts. Although long and longitudinal context is prominent in real-world patient and medical data, there is little work investigating LLM capabilities in such settings.
@@ 2.2 Retrieval-augmented Generation (RAG)
RAG has been leveraged in a prompt-based plug-in manner to address black-box long-context challenges (Yu et al., 2023). Current retrieval strategies are primarily: chunk-based (Izacard et al., 2021), index-based (Liu, 2022) and summarization-based (Sarthi et al., 2024). LLM-based large-scale retrievers for chunk-based approaches like BGE M3 (Chen et al., 2024a) give notable performance improvements. Furthermore, approaches like Self-RAG (Asai et al., 2024) outperform ChatGPT in open-domain QA through on-demand retrieval followed by generation and self-reflection.
Biomedical retrievers showed benefits in domain retrieval, with the MedCPT dense semantic retriever and re-ranker (Jin et al., 2023), achieving top performance on 6 biomedical tasks surpassing LLM-based counterparts. The BMRETRIEVER (Xu et al., 2024b) further surpasses MedCPT’s performance on most downstream tasks. More recently, methods like Med-RAG (Zhao et al., 2025) showed better diagnostics over EHRs, through enhancement by knowledge graph-elicited reasoning, while Self-BioRAG (Jeong et al., 2024) employs MedCPT and generates through self-reflection using a domain-specific LLM. Yet Fan et al. (2025b) underscore the medical retrieval challenges despite strong in-domain retrieval capabilities across large corpora, with models struggling with specialized medical content such as EHRs. Myers et al. (2025) further demonstrate variability in retrieval across EHR datasets. Therefore a thorough examination of RAG vs LC over realistic long-context patient-oriented QA is needed.
@@ 3 Approach
Since strong medical QA performance offers the potential of LLM integration in clinical workflows we aim to uncover LLM limitations given complex patient-oriented scenarios. We focus on content-comprehension settings using longitudinal EHRs, thus emphasizing relevant patient-centric long-context across time and documents.
@@
Our comprehensive approach spans across expert annotated datasets covering three task formulations: 1) MCQA, 2) Extractive, and 3) Open-ended QA. We assess each dataset across four settings with varying note relevance (see §4.2).
3.1 Dataset Selection Criteria
@@
The following criteria were used for benchmark selection in our study (full comparison in Table LABEL:tab:dataset_landscape):
@@
Beyond MCQA: Medical MCQA benchmarks i.e. PubMedQA (Jin et al., 2019), MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022) evaluate option selection instead of grounded synthesis and risk pretraining leakage from textbooks, which can inflate scores, providing little signal towards noisy clinical settings. Our evaluation goes beyond MCQA to both extractive and open-ended QA.
@@
Long-context Documents: To address LLM capabilities in handling real-world LC we seek benchmarks extending well-beyond 8K, with inclusion of patient-specific content.
Expert Annotated QA pairs grounded on multi-document EHRs were our choice, to avoid clinical reliability issues like non-expert curation (Fan et al., 2025a) or synthetic data (Adams et al., 2025).
Patient-centric Longitudinal Content: Prioritizing clinically-relevant QA with reasoning, we focus on MIMIC-based EHR datasets (Johnson et al., 2016, 2023), offering timestamped notes across document types with consistent format.
3.2 RAG Approach
@@
Retrievers are distinguished into sparse and dense. Sparse retrievers such as BM25 (Robertson and Zaragoza, 2009) and Splade (Formal et al., 2021) use overlapping terms to match queries with snippets, while dense retrievers encode them into embeddings and match them based on semantic similarity. Recently, using LLM-based embeddings for dense retrieval has become standard practice.
@@
Research on biomedical MCQA (Wang et al., 2024) and biomedical document retrieval (Luo et al., 2022) has shown that hybrid approaches combining sparse and dense retrieval outperform single component counterparts. Additionally, non-hybrid experiments on EHR subsets of varying granularity (Fan et al., 2025b) show no clear winner between sparse and dense retrievers. Finally, re-ranking of retrieved snippets also exhibits superior performance in biomedical applications (Jin et al., 2023; Wang et al., 2024; Sohn et al., 2025). Based on the above, we leverage a hybrid approach followed by re-ranking, presented in Figure 2. Clinical notes are segmented into 512-token chunks, forming the document collections $D=\{d_{1},d_{2},\dotsc,d_{n}\}$. The question respectively forms the query $Q$.
@@
Sparse retrieval: we employ a lexical retriever as a ranking function of each chunk $d_{j}$  [^1] Here terms document and chunk are used interchangeably.:
@@
Dense retrieval: we use the same encoder, $E$, for the query and the chunks and capture semantic similarity of such pairs using the cosine similarity:
@@
Reciprocal Rank Fusion (RRF): combines the top-$k$ chunks from each retriever accounting for each document’s rank, using a smoothing constant $k_{RRF}$:
@@
Late Interaction Reranking: measures how well every $Q$-token is semantically supported by at least one document token, to produce the final top-$k$ set:
@@
Document Ordering: we sort retrieved chunks temporally rather than by retrieval score due to its superior long-context performance (Yu et al., 2024).
@@ 3.3 Evaluation Methodology
Evaluating medical QA systems has traditionally relied on automatic metrics and MCQA benchmarks Jin et al. (2019, 2021); Pal et al. (2022), which probe memorized knowledge and underrepresent the complexity of clinical reasoning, and EHR-grounded context. Towards a clinically meaningful assessment, we adopt a multi-dimensional, reference-based scheme that complements lexical overlap with embedding-based semantic similarity and domain-adapted Natural Language Inference (NLI). In line with emerging clinical evaluation practices (e.g., MEDIC (Kanithi et al., 2024)) and the rise of LLM-as-a-judge, we employ a calibrated rubric and we prioritize widely available metrics while cross-checking signals for robustness. Our datasets comprise gold-standard QA pairs that are either expert-verified or expert-generated enabling a clinically centric evaluation:
METEOR Banerjee and Lavie (2005) to capture surface-level and lexical similarity.
BERTScore Zhang et al. (2019), computed with Clinical BioBERT embeddings Alsentzer et al. (2019) to assess semantic similarity.
NLI Scores, to capture the logical relationship between reference and candidate answers. We use a domain-adapted NLI model Deka et al. (2023) and then measure Factual Consistency using the probability of non-contradiction Song et al. (2024) and Factual Precision using the entailment probability.
LLM-as-a-judge, employed with a 5-point scale rubric on three aspects: Correctness capturing factual consistency with respect to the gold answer, penalizing contradictions while allowing compatible additional information, Completeness capturing recall by assessing the predicted answer’s coverage of information present in the reference, Faithfulness capturing precision assessing if the predicted answer only contains information that is supported by the reference.
Accuracy reported on the MCQA task.